a mover must put a circular mirror two meters in diameter through a one meter by 1.8 meter doorway. find the lenght of the diagonal of the doorway. will the mirror fit?

please help:)!

Answer by checkley77(12569)

1^2+1.7^2=x^2
1+3.24=x^2
x^2=4.24
x=sqrt4.24
x=2.06 meters.
Looks like a fit.